# Finetuning and Quantization of EEG-Based Foundational BioSignal Models on ECG and PPG Data for Blood Pressure Estimation

### v1

@@ III-A Architecture and Fine-Tuning
CEReBrO is then also available in three sizes—small (3.58M parameters), medium (39.95M parameters), and large (85.15M parameters).
@@
To measure the benefit of leveraging a pre-trained EEG encoder, we compare fine-tuning against training from scratch (i.e., random initialization). In each setting, we train models of three different sizes (small, medium, large) for 100 epochs and extend unfrozen-backbone runs to 200 epochs to assess longer-term convergence. We use Xavier Initialization [25] when training from scratch. Performance is evaluated on the MIMIC-III and VitalDB datasets in terms of MAE, SD, and coefficient of determination (), as well as clinical standards (British Hypertension Society (BHS) and Association for the Advancement of Medical Instrumentation (AAMI)).
@@ Acknowledgment
This project is supported by the Swiss National Science Foundation under the grant number 193813 (PEDESITE project) and by the ETH Future Computing Laboratory (EFCL), financed by a donation from Huawei Technologies.

### v2

@@ III-A Architecture and Fine-Tuning
CEReBrO is then also available in three sizes—small (3.58M parameters), medium (39.95M parameters), and large (85.15M parameters). During fine-tuning, we largely retained the original hyperparameters from the pre-trained CEREBRO model. We set the learning rate to 1e-4 and used a batch size of 1024, selected primarily to optimize GPU memory utilization rather than through an extensive hyperparameter search.
@@
To measure the benefit of leveraging a pre-trained EEG encoder, we compare fine-tuning against training from scratch (i.e., random initialization). In each setting, we train models of three different sizes (small, medium, large) for 100 epochs and extend unfrozen-backbone runs to 200 epochs to assess longer-term convergence. We use Xavier Initialization [25] when training from scratch. Performance is evaluated on the MIMIC-III and VitalDB datasets in terms of MAE, SD, and coefficient of determination (), as well as clinical standards (British Hypertension Society (BHS) and Association for the Advancement of Medical Instrumentation (AAMI)). To mitigate the risk of overfitting during the fine-tuning of these large models, we employed standard validation techniques. Data was partitioned into training, validation, and test sets, and we utilized an early stopping strategy based on performance monitoring on the validation set to select the best model checkpoint and prevent over-specialization to the training data Note that while the foundation model was pre-trained on EEG data, our BP estimation workflow uses only ECG and PPG signals during fine-tuning and inference; no patient-specific EEG calibration is required.
@@ Acknowledgment
This project is supported by the Swiss National Science Foundation under the grant number 193813 (PEDESITE project) and by the ETH Future Computing Laboratory (EFCL), financed by a donation from Huawei. This work was supported by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID lp12 on Alps.